What is KP at 1269°C for the reaction 2CO(g) + O2(g) ⇌ 2CO2(g) if Kc is 2.6 × 1014 at the same temperature? × 10 (Enter your answer in scientific notation.)
Answer #1

Answer

We know that the relation between Kp and Kc is Kp= Kc(RT)^delta n

here Kp= pressure equilibrium constant =?

Kc = concentration equilibrium cnstant = 2.6*10^14

R= gas constant = 0.0821 L.atm/mol.K

T= temperature = 1269+273.15= 1542.15 K  '

delta n= no of gaseous products - no of gaseous reactants = 2-(2+1)= -1

Kp= Kc*(RT)^delta n

Kc= 2.6*10^14 *(0.0821*1542.15)^-1 = 2.05*10^12